# Convergent Power Series for Anharmonic Chain with Periodic Forcing

### v3

@@ 5.1. Remark about non-uniqueness
This implies for the 00-site
@@ 7.1. Global stability of the periodic solution in the case of a harmonic chain on ${\mathbb{Z}}_{N}$
[…]
@@
where $\mathbb{F}(t)$ is a vector valued function, with $2(2N+1)$ components that are all 00, except the one corresponding to the momentum co-ordinate at $x=0$, where it equals ${\cal F}(t)$. In particular, ${\bf q}_{\rm p}(t)$ the $\theta$-periodic solution of (1.2) is given by
@@
It can be shown, see [30, Proposition 1.2] that there exists $C_{0}>0$ such that
@@ 7.2. Global stability via the approximation scheme
As a corollary of Theorem 7.3 we conclude that $\Big{(}{\bf q}\big{(}\cdot;s,\nu\big{)},{\bf p}\big{(}\cdot;s,\nu\big{)}\Big{)}$ converges to the periodic solution $\big{(}{\bf q}_{\rm p}(\cdot;\nu),{\bf p}_{\rm p}(\cdot;\nu)\big{)}$, as $s\to-\infty$, in $L^{2}_{\rm loc}[0,+\infty)$. Since the functions are bounded, together with their derivatives, the family is compact in the topology of uniform convergence on compact subsets of $[0,+\infty)$. The $L^{2}_{\rm loc}$ convergence allows us to identify the limit, as the periodic solution, which in particular implies the conclusion of Theorem 2.4.∎
@@ 7.2.2. Approximation scheme for the Laplace transform
00 0 $=\left[\Delta-\omega_{0}^{2}-\lambda^{2}-\gamma\lambda(\delta_{-N}(x)+\delta_{N}(x))\right]\widehat{q}_{x}^{(L)}(\lambda;s,\nu)$ (7.24)
@@
$\delta\widehat{v}_{x,L-1}(\lambda;s,\nu):=\widehat{v}_{x,L-1}(\lambda;s,\nu)-\widehat{v}_{x,L-1}^{({\rm p})}(\lambda;\nu).$
@@
If $\eta\leqslant-\sqrt{\omega_{0}^{2}-a}$, then we multiply both sides of (7.30) by $-1$ and (7.32) is still in force. Consequently choosing $\xi$ such that $2\xi\sqrt{\omega_{0}^{2}-a}\geqslant a$ we have that (7.31) holds for any $\eta$.
@@
Denote by $B_{R}$ the ball of radius $R>0$, centered at 00 in ${\mathbb{R}}^{2(2N+1)}$ and by ${\mathfrak{q}}(t)=\big{(}{\mathfrak{q}}_{x}(t)\big{)}$, $\dot{\mathfrak{q}}(t)=\big{(}\dot{\mathfrak{q}}_{x}(t)\big{)}$.
@@
We restore the dependence on the initial data $({\bf q},{\bf p})\in{\mathbb{R}}^{2(2N+1)}$ in the notation of $\delta q_{x}^{(0)}(t;s,{\bf q},{\bf p},\nu)$. From Lemmas 7.7 and 7.8 we immediately conclude the following.
@@ A.1. Green’s function of the operator $\lambda+\omega_{0}^{2}-\Delta$ on ${\mathbb{Z}}$
Suppose that $-\lambda\in\mathbb{C}\setminus{\cal I}$. The Green’s function of the operator $\lambda-\omega_{0}^{2}+\Delta$, where $\Delta$ is the free Laplacian on ${\mathbb{Z}}$ is given by
@@
considered for $|z|>1$. Then $z=\Phi_{+}(\zeta)$, $\zeta\not\in[-1,1]$. Condition $-\lambda\not\in{\cal I}$ is equivalent with $\zeta\not\in[-1,1]$. Furthermore $|\Phi_{+}\big{(}\chi(\lambda)\big{)}|>1$. We have then
@@ A.3. Green’s function for $\lambda+\omega^{2}_{0}+i\sigma(\delta_{-N}+\delta_{\ N})-\Delta$
For $x=-N$ and $x=N$ we get in particular the following system of equations on $H_{\lambda}^{(N)}(\pm N,y)$
@@ B.1. Some auxiliaries
It is an analytic function in $k\in\mathbb{T}$. Therefore, see e.g. [24, p. 27], there exist constants $A$ and $\rho>0$ such that
@@ Appendix D An application of the implicit function theorem in the proof of non-uniqueness of periodic solution
00 0 $=\big{[}(\omega m)^{2}-\omega_{0}^{2}+\Delta-i\gamma\omega m(\delta_{-N}+\delta_{N})\big{]}\widetilde{\overline{q}}_{x}(m;{\rm F})+{\rm F}\big{(}\delta_{1}(m)+\delta_{-1}(m)\big{)}\delta_{0}(x)$ (D.1)
@@
$\Phi\big{(}{\rm F},\widetilde{f}\big{)}_{x}(m):=\widetilde{f}_{x}(m)-H^{(N)}_{-(m\omega)^{2},\gamma m\omega}(x,0){\rm F}\big{(}\delta_{1}(m)+\delta_{-1}(m)\big{)}$ (D.4)
@@
$f_{x}(t):=\sum_{m\in{\mathbb{Z}}}\widetilde{f}_{x}(m)e^{im\omega t},\quad t\geqslant 0.$
@@
The Fréchet derivative of the right hand side of (D.3) with respect to the $q$ variable, at $(0,\widetilde{\overline{q}})$ is a linear operator ${\rm L}:{\ell}_{2,N}\to{\ell}_{2,N}$ given by
@@
It is easy to see that the operator ${\rm K}:{\ell}_{2,N}\to{\ell}_{2,N}$ is compact. We shall show that its null space is trivial. This would imply, see e.g. [27, Theorem 21.2.6, 238], that ${\rm L}$ is onto and its inverse is bounded. Using the implicit function theorem, see e.g. [10, Theorem 4.15.1], we conclude then that there exists $\overline{F}_{0}>0$ such that for any ${\rm F}\in(-\overline{F}_{0},\overline{F}_{0})$ equation (D.4) has a unique solution that gives rise to a $\theta$ periodic solution $\overline{q}_{{\rm p},x}(t;{\rm F})$ of (1.2) that satisfies
@@ Acknowledgements
J.L.L. thanks D. Huse and A. Dhar for useful discussions.

### v4

@@ 7.1. Proof of positivity of the work functional
This implies for the 0-site
@@
7.2. Proof of (3.22)
To conclude (3.22) it suffices to show that
$\lim_{N\to+\infty}\frac{1}{\theta}\int_{0}^{\theta}\big{[}p_{\pm N,{\rm p}}(t;\nu)\big{]}^{2}\mathrm{d}t=0.$ (7.1)
From Proposition 5.4 we have
$p_{\pm N,{\rm p}}(t;\nu)=\sum_{L=0}^{+\infty}\nu^{L}\dot{q}^{(L)}_{\pm N,{\rm p}}(t;\nu),$
where the series converges uniformly in $t$, as $N\to+\infty$, for $|\nu|<\nu_{0}$. Using this and estimate (5.27), to conclude (7.1) it suffices to prove that
$\lim_{N\to+\infty}\frac{1}{\theta}\int_{0}^{\theta}\big{[}q^{(L)}_{\pm N,{\rm p}}(t;\nu)\big{]}^{2}\mathrm{d}t=\lim_{N\to+\infty}\sum_{m\in{\mathbb{Z}}}|\widetilde{q}_{\pm N}^{(L)}(m)|^{2}=0$ (7.2)
for each $L=0,1,\ldots$. The latter is a consequence of formulas (5.15) and (5.16) and estimate (B.10).
@@ 8.1. Global stability of the periodic solution in the case of a harmonic chain on ${\mathbb{Z}}_{N}$
We denote by $\sigma(A)$ the set of all eigenvalues of $A$ and by ${\rm Re}\,\sigma(A)$ the set of all their real parts. Both here and in what follows by the norm of the matrix $M=[m_{i,j}]$ we understand its operator norm $\|M\|:=\sup_{\|x\|=1}\|Mx\|,$ where $\|\cdot\|$ is the euclidean norm of a vector.
@@
where $\mathbb{F}(t)$ is a vector valued function, with $2(2N+1)$ components that are all 0, except the one corresponding to the momentum co-ordinate at $x=0$, where it equals ${\cal F}(t)$. In particular, ${\bf q}_{\rm p}(t)$ the $\theta$-periodic solution of (2.3) is given by
@@
It can be shown, see [32, Proposition 1.2] that there exists $C_{0}>0$ such that
@@ 8.2. Global stability via the approximation scheme
As a corollary of Theorem 8.3 we conclude that $\Big{(}{\bf q}\big{(}\cdot;s,\nu\big{)},{\bf p}\big{(}\cdot;s,\nu\big{)}\Big{)}$ converges to the periodic solution $\big{(}{\bf q}_{\rm p}(\cdot;\nu),{\bf p}_{\rm p}(\cdot;\nu)\big{)}$, as $s\to$-$\infty$, in $L^{2}_{\rm loc}[0,+\infty)$. Since the functions are bounded, together with their derivatives, the family is compact in the topology of uniform convergence on compact subsets of $0,+\infty)$. The $L^{2}_{\rm loc}$ convergence allows us to identify the limit, as the periodic solution, which in particular implies the conclusion of Theorem 3.4.∎
@@ 8.2.2. Approximation scheme for the Laplace transform
$0$ $=\left[\Delta-\omega_{0}^{2}-\lambda^{2}-\gamma\lambda(\delta_{-N}(x)+\delta_{N}(x))\right]\widehat{q}_{x}^{(L)}(\lambda;s,\nu)$ (8.24)
@@
$\delta\widehat{v}_{x,L-1}(\lambda;s,\nu)=\widehat{v}_{x,L-1}(\lambda;s,\nu)-\widehat{v}_{x,L-1}^{({\rm p})}(\lambda;\nu).$
@@
If $\eta\leqslant-\sqrt{\omega_{0}^{2}-a}$, then we multiply both sides of (8.30) by 1-1 and (8.32) is still in force. Consequently choosing $\xi$ such that $2\xi\sqrt{\omega_{0}^{2}-a}\geqslant a$ we have that (8.31) holds for any $\eta$.
@@
Denote by $B_{R}$ the ball of radius $R>0$, centered at 0 in ${\mathbb{R}}^{2(2N+1)}$ and by ${\mathfrak{q}}(t)=\big{(}{\mathfrak{q}}_{x}(t)\big{)}$, $\dot{\mathfrak{q}}(t)=\big{(}\dot{\mathfrak{q}}_{x}(t)\big{)}$.
@@
We restore the dependence on the initial data $({\bf q},{\bf p})\in{\mathbb{R}}^{2(2N+1)}$ in the notation of $\delta q_{x}^{(0)}(t;s,{\bf q},{\bf p},\nu)$. From Lemmas 8.7 and 8.8 we immediately conclude the following.
@@ A.1. Green’s function of the operator $\lambda+\omega_{0}^{2}-\Delta$ on $\mathbb{Z}$}
Suppose that $\lambda\in\mathbb{C}\setminus{\cal I}$. The Green’s function of the operator $\lambda-\omega_{0}^{2}+\Delta$, where $\Delta$ is the free Laplacian on $\mathbb{Z}$} is given by
@@
considered for $|z|>1$. Then $z=\Phi_{+}(\zeta)$, $\zeta\not\in[-1,1]$. Condition $\lambda\not\in{\cal I}$ is equivalent with $\zeta\not\in[-1,1]$. Furthermore $|\Phi_{+}\big{(}\chi(\lambda)\big{)}|>1$. We have then
@@ A.3. Green’s function for $\lambda+\omega^{2}_{0}+i\sigma(\delta_{-N}+\delta_{\ N})-\Delta$
For x=Nx=-N and $x=N$ we get in particular the following system of equations on $H_{\lambda}^{(N)}(\pm N,y)$
@@ B.1. Some auxiliaries
It is an analytic function in $k\in\mathbb{T}$. Therefore, see e.g. [26, p. 27], there exist constants $A$ and $\rho>0$ such that
@@ Appendix D An application of the implicit function theorem in the proof of non-uniqueness of periodic solution
$0$ $=\big{[}(\omega m)^{2}-\omega_{0}^{2}+\Delta-i\gamma\omega m(\delta_{-N}+\delta_{N})\big{]}\widetilde{\overline{q}}_{x}(m;{\rm F})+{\rm F}\big{(}\delta_{1}(m)+\delta_{-1}(m)\big{)}\delta_{0}(x)$ (D.1)
@@
$\Phi\big{(}{\rm F},\widetilde{f}\big{)}_{x}(m)=\widetilde{f}_{x}(m)-H^{(N)}_{-(m\omega)^{2},\gamma m\omega}(x,0){\rm F}\big{(}\delta_{1}(m)+\delta_{-1}(m)\big{)}$ (D.4)
@@
$f_{x}(t)=\sum_{m\in{\mathbb{Z}}}\widetilde{f}_{x}(m)e^{im\omega t},\quad t\geqslant 0.$
@@
[…]
@@
It is easy to see that the operator ${\rm K}:{\ell}_{2,N}\to{\ell}_{2,N}$ is compact. We shall show that its null space is trivial. This would imply, see e.g. [29, Theorem 21.2.6, 238], that $\rm L$} is onto and its inverse is bounded. Using the implicit function theorem, see e.g. [10, Theorem 4.15.1], we conclude then that there exists $\overline{F}_{0}>0$ such that for any ${\rm F}\in(-\overline{F}_{0},\overline{F}_{0})$ equation (D.4) has a unique solution that gives rise to a $\theta$ periodic solution $\overline{q}_{{\rm p},x}(t;{\rm F})$ of (2.3) that satisfies
@@ Acknowledgements
We thank David Huse and Abhishek Dhar for useful discussions. J.L.L. and P.G. wish to express thanks to the Institute for Advanced Studies for its hospitality where part of this work was done.
Data availability
We do not analyse or generate any datasets, because our work proceeds within a theoretical and mathematical approach.
Conflict of interest statement
The authors declare no conflicts of interest regarding this manuscript.